OpenStudy (anonymous):

What is the equation of the line passing through (2, –3) and parallel to the line y = –6x – 1 in standard form?

OpenStudy (anonymous):

i keep getting 6x – y = –9 but i dont think that it is right

OpenStudy (anonymous):

well use y = m x + b m=-6 and you plug in coordinants

OpenStudy (anonymous):

-3=-6*2+b sfb b= 9 y=-6 x +9

OpenStudy (anonymous):

Since its parallel to the line then slope1=slope2 slope2 = -6. According to the law y2-y1=m(x2-x1) Use the given (2,-3) ^.^ (x1,y1) m=-6.. Substitute and form the equation.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Yw.